OpenStudy (anonymous):

please help!? <3 If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + y = 2 3x + 2y = 9 choices: 3(y + 2) + 2y = 9 3(–y + 2) + 2y = 9 3x + 2(y + 2) = 9 3x + 2(–y + 2) = 9 I think it might be C...but I'n most likley wrong.

OpenStudy (anonymous):

Here this might help you :) http://openstudy.com/study#/updates/4ecfb176e4b04e045af38f49

OpenStudy (aravindg):

ok can you first tell me what x and y are from first equation?

OpenStudy (anonymous):

uhm... 1?

OpenStudy (aravindg):

x + y = 2 write x in terms of y and y in terms of x

OpenStudy (aravindg):

x=? y=?

OpenStudy (anonymous):

3(2-y)+2y = 9 x + y = 2 x = 2 - y 3(2-y) + 2y = 9

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

let x=(2-y) 3x + 2y = 9 .............eqn 2 put d value of x in eqn 2 3(2-y) + 2y=9 6-3y+2y=9 6-y=9 y=-3

OpenStudy (anonymous):

so avin y as -3 in eqn 2 3x+2(-3)=9 3x-6=9 3x=15 x=5